OpenStudy (anonymous):

What is the range of this data set? 319, 127, 183, 248, 655, 943, 891, 781 535 687 816 926

OpenStudy (ja1):

Range is the smallest number - the largest so...

OpenStudy (anonymous):

943-127= 816

OpenStudy (ja1):

127 - 943= ? (flip the equation)

OpenStudy (anonymous):

816

OpenStudy (ja1):

Well yes that's it @k.rajabhishek

OpenStudy (goldphenoix):

To find the range, you must do the maximum number subracted by the minimum number. The maximum number I see is 943. The minimum number I see is 127. Now do: 943 - 127 = 816. I hope this help! :)